solve by the substitution method 3x+7y=39 and -8+y=73
Answer by fcabanski(1391)

-8+y=73


Add 8 to both sides.


y=81.


Substitute that into the first equation.


3x+7(81)=39


3x+567=39


3x=-528


x=-176

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com